# Real-Time Sampling-Based Safe Motion Planning for Robotic Manipulators in Dynamic Environments

### v1

@@ 53                   $replanning\leftarrow\textnormal{{{true}}}$
First, the predefined path $\boldsymbol{Q}_{pred}$ can be obtained by any static planner (e.g., RRT-Connect [10], RGBT-Connect [51], or RGBMT* [55]). At the beginning, $\boldsymbol{Q}_{trav}$ only contains $\boldsymbol{q}_{start}$. A local variable $replanning$ denotes whether the replanning is required, and it is initially set to false. Moreover, there is a local variable $status$ indicating the robot’s current status (reached, advanced, trapped – inherited from RRT [6]).
@@
Lines 11 are being executed, until $\boldsymbol{q}_{curr}$ reaches $\boldsymbol{q}_{goal}$ or the collision occurs. At the beginning, the function $\mathtt{generateHorizon}$ computes a horizon $\boldsymbol{Q}_{h}$ using $\boldsymbol{q}_{curr}$, $\boldsymbol{Q}_{pred}$, and $status$. For the computation of generalized bur, we use the generalization of $d_{c}$, as proposed in [56]. For convenience, we utilize constant values $d_{i}$ for $i$-th separate robot’s link, $i\in\{1,2,\dots,n\}$, which are acquired by the $\mathtt{computeDistances}$ function within a vector $\boldsymbol{d}=[d_{1},d_{2},\dots,$ $d_{n}]^{T}$. Such piecewise constant function can be easily obtained as a byproduct of computing the minimal distance $d_{c}=\min\{\boldsymbol{d}\}$, since distances for each pair link-obstacle $d_{i,j}$, $j\in\{1,2,\dots,N_{obs}\}$, are computed separately anyway. Therefore, $d_{i}=\min\{d_{i,1},d_{i,2},\dots,d_{i,N_{obs}}\}$.
@@
Afterwards, $\boldsymbol{Q}_{h}$ is updated within the $\mathtt{updateHorizon}$ routine, including the update of $N_{h}$ according to (1). For the generation of random and lateral spines, $\boldsymbol{q}_{next}$ is required. In case it is not determined, the first node from $\boldsymbol{Q}_{h}$ is chosen as $\boldsymbol{q}_{next}$. A generalized bur $GBur$, starting in $\boldsymbol{q}_{curr}$ and oriented towards all nodes from $\boldsymbol{Q}_{h}$, is generated by the function $\mathtt{generateGBur}$. Thereafter, the $\mathtt{computeNodeWeights}$ routine computes all bur node weights $\boldsymbol{w}$. The $\mathtt{getNextState}$ function returns $\boldsymbol{q}_{next}$ based on previously obtained weights. In case all $\boldsymbol{Q}_{h}$ nodes are critical ones, $status$ becomes trapped, suggesting that $\boldsymbol{q}_{next}$ equals $\boldsymbol{q}_{curr}$. Consequently, the robot will eventually stop (which will be elaborated in detail in Sec. IV), and $replanning$ is set to true to trigger the search for better nodes.
@@
Since the robot should advance towards $\boldsymbol{q}_{next}$ by reaching a new configuration $\boldsymbol{q}_{new}$, a time parameterization of a local path $\overline{\boldsymbol{q}_{curr}\boldsymbol{q}_{next}}$ is required regarding $\mathcal{K}$. As a result, a trajectory $\boldsymbol{\pi}_{curr}(t)=\boldsymbol{\pi}[\boldsymbol{q}_{curr},\boldsymbol{q}_{next}]$ containing $\boldsymbol{q}_{new}$ is obtained, thus $\boldsymbol{q}_{curr}$ is updated to $\boldsymbol{q}_{new}$ extending $\boldsymbol{Q}_{trav}$ by line 1. This procedure is computed within the $\mathtt{updateCurrState}$ function depending on the input parameter $safe\_on$, which indicates whether the safe variant of DRGBT, so-called DRGBT-safe (see Sec. IV) is used for computing $\boldsymbol{\pi}_{curr}(t)$. The status of this motion is tracked via $status$. When it becomes different from advanced, line 1 is executed again. If $status$ is reached, it means that $\boldsymbol{q}_{next}$ is reached, thus $\boldsymbol{Q}_{h}$ will be updated with the new nodes from $\boldsymbol{Q}_{pred}$. Otherwise, when $status$ is trapped, $\boldsymbol{Q}_{h}$ needs to be generated using only random nodes.
@@
Depending on the variable $replanning$ and the function $\mathtt{whetherToReplan}$, it is decided whether the replanning occurs. Using any static planner, the path is replanned from $\boldsymbol{q}_{curr}$ to $\boldsymbol{q}_{goal}$. When a new path $\boldsymbol{Q}_{pred,new}$ is found, the function $\mathtt{updatePath}$ replaces the previous one, and $replanning$ is set to false. Moreover, $status$ becomes reached in order to generate a new horizon $\boldsymbol{Q}_{h}$ by line 1. Conversely, if the new path cannot be found, $replanning$ is set to true, thus its replanning is explicitly required during the next iteration. It is worth mentioning that $\mathtt{updatePath}$ also refines $\boldsymbol{Q}_{pred,new}$ by modifying its all nodes such that the (weighted) Euclidean distance between each two adjacent nodes is not greater than $\|\boldsymbol{\omega}_{max}\|\cdot T$, which represents maximal distance the robot can cover in $\mathcal{C}$-space during an iteration time $T$. The refined path remains the same geometrically, yet only reallocates its nodes to the carefully chosen locations. It turned out that such pre-processing yields better performance of the overall algorithm.
@@ III Scheduling Framework of the Algorithm
The top priority for real-time systems is that the used scheduling algorithm produces a predictable schedule, i.e., the next executable task must be known at each instance of time. This section is mainly concerned with the execution times of individual tasks within the algorithm. Extensive simulation study reveals the main challenges for the real-time execution, but also the appropriate mechanism to address them. Commonly used approaches to real-time scheduling are clock-driven, priority-driven, and round-robin [57]. Recently, priority-driven methods, such as earliest deadline first (EDF), rate-monotonic scheduling (RMS), fixed priority scheduling (FPS), dynamic priority scheduling (DPS), and their variants have become very popular. They are applied in many real-time operating systems due to their simplicity and predictability. Following a number of considerations, EDF method turned out to be the most convenient one, thus it will be applied hereafter. The most of theoretical analysis stems from [57, 58, 59] and [60]. To accommodate the algorithm for schedulability analysis, it is decomposed into two main tasks: Task 1 (T1) – computing the robot’s next configuration, and Task 2 (T2) – replanning the predefined path.
Figure 4: EDF scheduling of T1 and T2 with two priority levels
@@ III-A Task 1: Computing the Robot’s Next Configuration
T1 is periodic task with a period $T$ (the same as the algorithm iteration time) and an execution/maximal time $e_{1}$. It is released/activated in time instances $t=kT$, $k\in\mathbb{N}_{0}$, as indicated by Fig. 4. The task includes lines 11 from Alg. 1. Since we want the robot to avoid collision with obstacles, it needs to receive commands generated by this task timely, i.e., each $T$. Therefore, T1 must be categorized as hard real-time task with its relative deadline of $D_{1}=T$.
It should be mentioned that the new desired configuration $\boldsymbol{q}_{new}\in\boldsymbol{\pi}[\boldsymbol{q}_{curr},\boldsymbol{q}_{next}]$ is carefully computed within the $\mathtt{updateCurrState}$ function in order to make the robot capable of reaching $\boldsymbol{q}_{new}$ for the duration of $T$. Whether the trajectory $\boldsymbol{\pi}[\boldsymbol{q}_{curr},\boldsymbol{q}_{new}]$ will remain collision-free during the current iteration will be discussed in Sec. IV.
@@ III-B Task 2: Replanning the Predefined Path
This task is sporadic since it is conditioned by both the variable $replanning$ and the function $\mathtt{whetherToReplan}$. It is subjected to a sporadic server, which is de facto periodic as T1 (with the same period $T$). T2 is released at the same time instances as T1, i.e., in $t=kT$, and its relative deadline is $D_{2}$, which is equal to $D_{1}=T$. Due to the same deadlines, $D_{1}$ and $D_{2}$, both tasks have the same priority, but the EDF scheduler lets T1 be executed always first giving it a top-level priority. Immediately after its completion, the server checks (by line 1 from Alg. 1) whether T2 needs to be executed. If yes, lines 11 must be executed within its execution time
$e_{2}=T-e_{1},\vspace{-0.2cm}$ (2)
which stems from its dependency on $e_{1}$. More precisely, T2 executes until it completes or its time slice $e_{2}$ expires, as indicated by the clock interrupt in Fig. 4 by red lines. If T2 does not complete, it is suspended when T1 is released (in $t=kT$), since it has been assigned a lower priority compared to T1. In this case, the new path is not found, and the replanning is explicitly required in the subsequent iterations. Nevertheless, we assume T2 is successfully executed in terms of scheduling theory. The choice of such task priority levels will be elaborated in more detail later (Subsec. V-C).
T2 is categorized as hard real-time task regarding two aspects. First, if replanning is required, it implies that the actual predefined path is not necessarily collision-free, and the robot is possibly in an emergency requiring the new path as soon as possible. Second, if the replanning breaks its deadline, the following task T1 will delay its execution, i.e., it will not be released in $t=kT$. Accordingly, if T2 does not complete successfully (when the new predefined path is not found within $e_{2}$ although it can in fact exist), the robot may exploit the previously predefined path. Moreover, adaptive horizon, which mimics the set of range sensors in $\mathcal{C}$-space, will locally control the robot to avoid collision. The details on safety guarantees will be discussed in Sec. IV.
@@
To ensure that both tasks meet their deadlines in hard real-time, the following inequality must hold [57]:
$\frac{e_{1}}{D_{1}}+\frac{e_{2}}{D_{2}}\leq 1\implies\frac{e_{1}}{T}+\frac{e_{2}}{T}\leq 1,\vspace{-0.15cm}$ (3)
where $u_{i}=\frac{e_{i}}{T}$ determines a utilization of Task $i$, $i\in\{1,2\}$. After substituting (2) into (3), it yields $1\leq 1$, which proves the correctness of our approach.
@@
IV Guaranteed Safe Motion of the Robot in Dynamic Environments
@@
In case of human presence in the robot’s vicinity, it is generally undesirable to unconditionally aim at fast, minimum-time trajectories. Thus, we upgrade our approach (for generating trajectories within the $\mathtt{updateCurrState}$ routine from Alg. 1) in terms of automatically scaling the robot’s velocity in order to guarantee a safe motion of the robot. The key factor for this upgrade is the information about the minimal robot-obstacle distance $d_{c}$. In particular, the robot should be able to controllably decelerate to zero speed without causing a collision with the environment [^1] This feature is consistent with the stop category 2 defined in the standard ISO 10218 Robots and robotic devices – Safety requirements for industrial robots. Our approach is also motivated by the speed and separation monitoring (SSM) concept (described in the standard ISO/TS 15066 Robots and robotic devices – Collaborative robots).. The collision is allowed to occur once the robot has stopped, which will be referred to as the type 2 collision, differing from the type 1 collision that assumes the robot-obstacle contact with the robot having non-zero speed.
@@ IV-A Dynamic Expanded Bubbles and Burs
As mentioned in Subsec. II-D, we use the piecewise constant approximation of the profile $d(s)$ via constant values $d_{i}$ for $i$-th separate robot’s link, which are represented by the vector $\boldsymbol{d}=[d_{1},d_{2},\dots,d_{n}]^{T}$, as proposed in [56]. Furthermore, the points $\boldsymbol{R}_{i,j}$ and $\boldsymbol{O}_{i,j}$ are obtained as a byproduct of computing $d_{c}$, which determine the nearest points between $i$-th robot’s link and $\mathcal{WO}_{j}$, respectively. These points uniquely define the plane $\mathcal{P}_{i,j}$ (e.g., $\mathcal{P}_{i,j}(t)$ for $t=t_{0}$ from Fig. 7 (b)) that conveniently divides the workspace $\mathcal{W}$ into two halfspaces, free one containing $i$-th robot’s link, and the one occupied by $\mathcal{WO}_{j}$, as proven in [51]. All the planes (corresponding to link-obstacle pairs) can be expressed analytically and their parameters conveniently structured within a matrix $\boldsymbol{\mathcal{P}}$.
All captured obstacles are allowed to move in any direction during the current iteration as long as their velocities do not exceed $v_{obs}$. Theoretically, it is allowed for the new obstacles to suddenly emerge during the current iteration in a region $\mathcal{W}_{occ}\subset\mathcal{W}$ obtained as a union of the occupied halfspaces w.r.t. each $\mathcal{P}_{i,j}(t)$. The sudden appearance of new obstacles in the safe region $\mathcal{W}_{safe}=\mathcal{W}\backslash\mathcal{W}_{occ}$ is not allowed.
@@
$\mathcal{DEB}(\widetilde{\boldsymbol{q}},\boldsymbol{d},v_{obs})=\left\{\boldsymbol{y}(\widetilde{\boldsymbol{q}},t)~{}\Big{|}~{}\overline{\rho}_{i}(\widetilde{\boldsymbol{q}},t)+v_{obs}t\leq d_{i}\right\},\vspace{-0.2cm}$ (4)
for $i\in\{1,2,\dots,n\}$ and $t\in\left[0,t^{*}\right]$, where $\overline{\rho}_{i}(\widetilde{\boldsymbol{q}},t^{*})+v_{obs}t^{*}=d_{i}$, $\forall i$. The sum $\overline{\rho}_{i}(\widetilde{\boldsymbol{q}},t)=\sum_{j=1}^{i}r_{i,j}\left|y_{i}(\widetilde{\boldsymbol{q}},t)-q_{i}\right|$ determines a conservative upper bound on the displacement of any point on the manipulator $\mathcal{R}_{i}$ when changing its configuration from $\boldsymbol{q}$ to an arbitrary configuration $\boldsymbol{y}(\widetilde{\boldsymbol{q}},t)$ during the time $t$. The weight $r_{i,j}$ represents a radius of the cylinder coaxial with $j$-th joint enclosing all parts of the robot $\mathcal{R}_{i}$ starting from $j$-th joint when the robot is at the configuration $\boldsymbol{q}$. The term $\mathcal{R}_{i}$ designates the “incomplete” robot starting at its base and ending up at its $i$-th joint.
Generally, any motion within $\mathcal{DEB}$ means that no point on the robot $\mathcal{R}_{i}$ will move more than $d_{i}-v_{obs}t$ for a given time $t$, thus no collision can occur. Such a defined bubble differs from the one from [56] in terms of reducing the initial value of each $d_{i}$ for $v_{obs}t$. The rationale is that each $\mathcal{WO}_{j}$, may cover the distance of $v_{obs}t$ in the worst-case scenario, i.e., it may approach the robot’s $i$-th link at its highest speed $v_{obs}$. This observation is equivalent to considering the plane $\mathcal{P}_{i,j}$ which artificially enlarges $\mathcal{WO}_{j}$ by following a velocity vector $\boldsymbol{v}_{\mathcal{P}_{i,j}}=\frac{\boldsymbol{R}_{i,j}-\boldsymbol{O}_{i,j}}{\|\boldsymbol{R}_{i,j}-\boldsymbol{O}_{i,j}\|}v_{obs}$ (as depicted in Fig. 7 (b)–(d)).
Fig. 5 illustrates the concept of dynamic bubbles. The workspace of a planar 2-DoF robot in four different configurations with two box-shaped obstacles is shown in Fig. 5 (left). Changing $v_{obs}$ from 00 to $6\,\mathrm{[\frac{m}{s}]}$ affects the $\mathcal{DEB}$ with a root in $\boldsymbol{q}_{0}=\left[-\frac{\pi}{4},0\right]^{T}$ (blue configuration) is revealed by Fig. 5 (right). Note that the case for $v_{obs}=0$ corresponds with the expanded bubble from [56]. For higher velocities, $\mathcal{DEB}$s get smaller, and their shape deviates from the diamond-like one. The initial velocity and acceleration used for time parameterization of trajectories are set to zero in all cases.

### v2

@@ 53        $replanning\leftarrow\textnormal{{{true}}}$
First, the predefined path $\boldsymbol{Q}_{pred}$ can be obtained by any static planner (e.g., RRT-Connect [10], RGBT-Connect [51], or RGBMT* [56]). At the beginning, $\boldsymbol{Q}_{trav}$ only contains $\boldsymbol{q}_{start}$. A local variable replanningreplanning denotes whether the replanning is required, and it is initially set to false. Moreover, there is a local variable statusstatus indicating the robot’s current status (reached, advanced, trapped – inherited from RRT [6]).
@@
Lines 11 are being executed, until $\boldsymbol{q}_{curr}$ reaches $\boldsymbol{q}_{goal}$ or the collision occurs. At the beginning, the function $\mathtt{generateHorizon}$ computes a horizon $\boldsymbol{Q}_{h}$ using $\boldsymbol{q}_{curr}$, $\boldsymbol{Q}_{pred}$, and statusstatus. For the computation of generalized bur, we use the generalization of $d_{c}$, as proposed in [57]. For convenience, we utilize constant values $d_{i}$ for $i$-th separate robot’s link, $i\in\{1,2,\dots,n\}$, which are acquired by the $\mathtt{computeDistances}$ function within a vector $\boldsymbol{d}=[d_{1},d_{2},\dots,$ $d_{n}]^{T}$. Such piecewise constant function can be easily obtained as a byproduct of computing the minimal distance $d_{c}=\min\{\boldsymbol{d}\}$, since distances for each pair link-obstacle $d_{i,j}$, $j\in\{1,2,\dots,N_{obs}\}$, are computed separately anyway. Therefore, $d_{i}=\min\{d_{i,1},d_{i,2},\dots,d_{i,N_{obs}}\}$.
@@
Afterward, $\boldsymbol{Q}_{h}$ is updated within the $\mathtt{updateHorizon}$ routine, including the update of $N_{h}$ according to (1). For the generation of random and lateral spines, $\boldsymbol{q}_{next}$ is required. In case it is not determined, the first node from $\boldsymbol{Q}_{h}$ is chosen as $\boldsymbol{q}_{next}$. A generalized bur $GBur$, starting in $\boldsymbol{q}_{curr}$ and oriented towards all nodes from $\boldsymbol{Q}_{h}$, is generated by the function $\mathtt{generateGBur}$. Thereafter, the $\mathtt{computeNodeWeights}$ routine computes all bur node weights $\boldsymbol{w}$. The $\mathtt{getNextState}$ function returns $\boldsymbol{q}_{next}$ based on previously obtained weights. In case all $\boldsymbol{Q}_{h}$ nodes are critical ones, statusstatus becomes trapped, suggesting that $\boldsymbol{q}_{next}$ equals $\boldsymbol{q}_{curr}$. Consequently, the robot will eventually stop (which will be elaborated in detail in Sec. IV), and replanningreplanning is set to true to trigger the search for better nodes.
@@
Since the robot should advance towards $\boldsymbol{q}_{next}$ by reaching a new configuration $\boldsymbol{q}_{new}$, a time parameterization of a local path $\overline{\boldsymbol{q}_{curr}\boldsymbol{q}_{next}}$ is required regarding $\mathcal{K}$. As a result, a trajectory $\boldsymbol{\pi}_{curr}(t)=\boldsymbol{\pi}[\boldsymbol{q}_{curr},\boldsymbol{q}_{next}]$ containing $\boldsymbol{q}_{new}$ is obtained, thus $\boldsymbol{q}_{curr}$ is updated to $\boldsymbol{q}_{new}$ extending $\boldsymbol{Q}_{trav}$ by line 1. This procedure is computed within the $\mathtt{updateCurrState}$ function depending on the input parameter safe_onsafe\_on, which indicates whether the safe variant of DRGBT, so-called DRGBT-safe (see Sec. IV) is used for computing $\boldsymbol{\pi}_{curr}(t)$. The status of this motion is tracked via statusstatus. When it becomes different from advanced, line 1 is executed again. If statusstatus is reached, it means that $\boldsymbol{q}_{next}$ is reached, thus $\boldsymbol{Q}_{h}$ will be updated with the new nodes from $\boldsymbol{Q}_{pred}$. Otherwise, when statusstatus is trapped, $\boldsymbol{Q}_{h}$ needs to be generated using only random nodes.
@@
Depending on the variable replanningreplanning and the function $\mathtt{whetherToReplan}$, it is decided whether the replanning occurs. Using any static planner, the path is replanned from $\boldsymbol{q}_{curr}$ to $\boldsymbol{q}_{goal}$. When a new path $\boldsymbol{Q}_{pred,new}$ is found, the function $\mathtt{updatePath}$ replaces the previous one, and replanningreplanning is set to false. Moreover, statusstatus becomes reached in order to generate a new horizon $\boldsymbol{Q}_{h}$ by line 1. Conversely, if the new path cannot be found, replanningreplanning is set to true, thus its replanning is explicitly required during the next iteration. It is worth mentioning that $\mathtt{updatePath}$ also refines $\boldsymbol{Q}_{pred,new}$ by modifying its all nodes such that the (weighted) Euclidean distance between each two adjacent nodes is not greater than $\|\boldsymbol{\omega}_{max}\|\cdot T$, which represents maximal distance the robot can cover in $\mathcal{C}$-space during an iteration time $T$. The refined path remains the same geometrically, yet only reallocates its nodes to the carefully chosen locations. It turned out that such pre-processing yields better performance of the overall algorithm.
@@ III Scheduling Framework of the Algorithm
The results from our prior work [50] clearly indicate that the original form of DRGBT is unsuitable for real-time execution, as its iteration time $T$ fluctuates significantly and heavily depends on the scenario. Moreover, $T$ might be infinite theoretically (e.g., if the replanner cannot find a solution at all). This variability motivated us to introduce a proper scheduling, supported by schedulability analysis and corresponding conditions for hard real-time operation. If we assume that the environment is perceived with the frequency $f_{perc}=1/T_{perc}$, then it is possible to miss new perception measurements in case $T>T_{perc}$. Yet, such measurements could be crucial for the robot to avoid collision. Therefore, this calls for introducing proper scheduling which limits $T$, i.e., some scheduling technique must be applied to the algorithm.
A key requirement of real-time systems is that the used scheduler produces a fully predictable task execution sequence at each time instance. Accordingly, this section is mainly concerned with how DRGBT’s tasks can be properly handled by a scheduler. Commonly used approaches to real-time scheduling follows clock-driven, priority-driven, and round-robin schemes [58]. Recently, priority-driven methods, such as earliest deadline first (EDF), rate-monotonic scheduling (RMS), fixed priority scheduling (FPS), dynamic priority scheduling (DPS), and their variants have been particularly popular. They are applied in many real-time operating systems due to their simplicity and predictability. Considering a wide variety of available scheduling algorithms and the corresponding theoretical analysis of schedulability ([58, 59, 60] and [61]), EDF method turned out to be the most convenient and will be applied thereafter.
Consider $N$ periodic tasks with periods $T_{i}$, relative deadlines $D_{i}$, and a worst-case/maximum execution times $e_{i}$, for $i\in\{1,2,\dots,N\}$. Under EDF scheduling, assuming all tasks are independent and preemptable, sufficient condition which guarantees that all tasks will meet their deadlines on a single-core processor is given as (see [58] for proof)
$\sum_{i=1}^{N}\frac{e_{i}}{\min\{D_{i},T_{i}\}}\leq 1.\vskip-5.69046pt$ (2)
This inequality guides the selection of task timing constraints. To accommodate the algorithm for schedulability analysis, it is decomposed into two main tasks described in the sequel.
Figure 4: EDF scheduling of T1 and T2 with two priority levels.
@@ III-A Task 1: Computing the Robot’s Next Configuration
Task 1 (T1) is defined as a strictly periodic hard real‐time task with a period $T$ (identical to the algorithm iteration time) and worst‐case/maximum execution time $e_{1}$. T1 is released/activated at $t=kT$, $k\in\mathbb{N}_{0}$, as depicted in Fig. 4. Upon each release, T1 deterministically executes the sequence of subtasks corresponding to lines 11 from Alg. 1. To ensure continuous collision avoidance, T1 must issue control commands at a maximum interval of $T$, thereby imposing a relative deadline $D_{1}=T$.
The new target configuration $\boldsymbol{q}_{new}\in\boldsymbol{\pi}[\boldsymbol{q}_{curr},\boldsymbol{q}_{next}]$ is carefully computed via $\mathtt{updateCurrState}$ routine to guarantee reachability of $\boldsymbol{q}_{new}$ within $T$. The collision‐free property of the resulting trajectory $\boldsymbol{\pi}[\boldsymbol{q}_{curr},\boldsymbol{q}_{new}]$ over each iteration is examined in Sec. IV.
@@ III-B Task 2: Replanning the Predefined Path
Task 2 (T2) is characterized as a sporadic task since it is triggered by both the variable replanningreplanning and the function $\mathtt{whetherToReplan}$ (line 1 in Alg. 1). It is wrapped in a sporadic server, which operates periodically in synchrony with T1 (with the same period $T$). T2 has the same activation schedule as T1, i.e., in $t=kT$, and its relative deadline is $D_{2}$, which is equal to $D_{1}=T$. Due to the same release times and same relative deadlines, jobs (instances) of both tasks share the same priority since the sequences of the absolute deadlines for their jobs (instances) are identical. However, EDF scheduler lets T1 be executed always first giving it a top-level priority. Post-T1, the server checks line 1 from Alg. 1, and, if needed, T2 must be executed by lines 11 during the time
$e_{2}=T-e_{1},\vskip-5.69046pt$ (3)
which stems from its dependency on $e_{1}$. More precisely, T2 executes either until completion or until its allocated time slice $e_{2}$ expires, as illustrated by the clock interrupts (red lines) in Fig. 4. If T2 does not complete by its deadline, it is forcibly terminated at the subsequent release of T1 (i.e., at $t=kT$). Such termination guarantees that T2 never exceeds its assigned execution budget $e_{2}$. If T2 is terminated, the new path is not found, and replanning must be explicitly re-invoked in subsequent iterations. Nevertheless, from a scheduling perspective, T2 is correctly completed. The rationale behind this priority assignment is further detailed in Subsec. V-C.
T2 is classified as a hard real-time task for two critical reasons. First, a replanning request indicates a potential emergency situation requiring an immediate new path to maintain safety. Second, if T2 misses deadline, it delays the release of T1. Accordingly, in cases where T2 fails to generate a new path within $e_{2}$ (though a feasible solution might exist), the robot may rely on the previous path. Additionally, the adaptive horizon, which emulates range sensing in $\mathcal{C}$-space, provides local collision avoidance and can ultimately initiate an emergency stopping (see further details in Sec. IV).
@@
To ensure both tasks meet their deadlines in hard real-time, the sufficient condition (2), for $D_{1}=D_{2}=T$, simplifies to
$\frac{e_{1}}{\min\{D_{1},T\}}+\frac{e_{2}}{\min\{D_{2},T\}}\leq 1\implies\frac{e_{1}}{T}+\frac{e_{2}}{T}\leq 1,\vskip-4.26773pt$ (4)
where $u_{i}=\frac{e_{i}}{T}$ denotes the utilization of Task $i$. Substituting (3) into (4) yields $1\leq 1$, proving schedulability for this setup.
It is important to emphasize that the schedulability analysis is primarily focused on T1, as execution times of its subtasks are accurately characterized and predictable under worst-case conditions (as it will be revealed by Figs. 11 and 12 in Sec. V). Once the deadline for T1 is rigorously satisfied, the temporal constraints for T2 arise as a consequence by (3).
@@
IV Safe Motion of the Robot in Dynamic Environments Under Bounded Obstacle Velocity
@@
In case of human presence in the robot’s vicinity, it is generally undesirable to unconditionally aim at fast, minimum-time trajectories. Thus, we upgrade our approach (for generating trajectories within the $\mathtt{updateCurrState}$ routine from Alg. 1) in terms of automatically scaling the robot’s velocity in order to guarantee a safe motion of the robot. The key factor for this upgrade is the information about the minimal robot-obstacle distance $d_{c}$. In particular, the robot should be able to controllably decelerate to zero speed without causing a collision with the environment [^1] This feature is consistent with the stop category 2 defined in the standard ISO 10218 Robots and robotic devices – Safety requirements for industrial robots. Our approach is also motivated by the speed and separation monitoring (SSM) concept (described in the standard ISO/TS 15066 Robots and robotic devices – Collaborative robots).. The collision is allowed to occur once the robot has stopped, which will be referred to as the type II collision, differing from the type I collision that assumes the robot-obstacle contact with the robot having non-zero speed.
@@ IV-A Dynamic Expanded Bubbles and Burs
As mentioned in Subsec. II-E, we use the piecewise constant approximation of the profile $d(s)$ via constant values $d_{i}$ for $i$-th separate robot’s link, which are represented by the vector $\boldsymbol{d}=[d_{1},d_{2},\dots,d_{n}]^{T}$, as proposed in [57]. Furthermore, the points $\boldsymbol{R}_{i,j}$ and $\boldsymbol{O}_{i,j}$ are obtained as a byproduct of computing $d_{c}$, which determine the nearest points between $i$-th robot’s link and $\mathcal{WO}_{j}$, respectively. These points uniquely define the plane $\mathcal{P}_{i,j}$ (e.g., $\mathcal{P}_{i,j}(t)$ for $t=t_{0}$ from Fig. 7 (b)) that conveniently divides the workspace $\mathcal{W}$ into two halfspaces, free one containing $i$-th robot’s link, and the one occupied by $\mathcal{WO}_{j}$, as proven in [51]. All the planes (corresponding to link-obstacle pairs) can be expressed analytically and their parameters conveniently structured within a matrix $\boldsymbol{\mathcal{P}}$.
It is worth stressing that the algorithm assumes that obstacles can move unpredictably in any direction during each iteration as long as the magnitude of their velocities does not exceed a maximal obstacle velocity $v_{obs}$. We do not assume any bounds on obstacles’ acceleration, thus, in theory, it can be infinite. Theoretically, it is allowed for the new obstacles to suddenly emerge during the current iteration in a region $\mathcal{W}_{occ}\subset\mathcal{W}$ obtained as a union of the occupied halfspaces w.r.t. each $\mathcal{P}_{i,j}(t)$. The sudden appearance of new obstacles in the safe region $\mathcal{W}_{safe}=\mathcal{W}\backslash\mathcal{W}_{occ}$ is not allowed.
@@
$\mathcal{DEB}(\widetilde{\boldsymbol{q}},\boldsymbol{d},v_{obs})=\left\{\boldsymbol{y}(\widetilde{\boldsymbol{q}},t)~\Big{|}~\overline{\rho}_{i}(\widetilde{\boldsymbol{q}},t)+v_{obs}t\leq d_{i}\right\},\vskip-5.69046pt$ (5)
for $i\in\{1,2,\dots,n\}$ and $t\in\left[0,t^{*}\right]$, where $\overline{\rho}_{i}(\widetilde{\boldsymbol{q}},t^{*})+v_{obs}t^{*}=d_{i}$, $\forall i$. The sum $\overline{\rho}_{i}(\widetilde{\boldsymbol{q}},t)=\sum_{j=1}^{i}r_{i,j}\left|y_{j}(\widetilde{\boldsymbol{q}},t)-q_{j}\right|$ determines a conservative upper bound on the displacement of any point on the manipulator $\mathcal{R}_{i}$ when changing its configuration from $\boldsymbol{q}$ to an arbitrary configuration $\boldsymbol{y}(\widetilde{\boldsymbol{q}},t)$ during the time $t$. The weight $r_{i,j}$ represents a radius of the cylinder coaxial with $j$-th joint enclosing all parts of the robot $\mathcal{R}_{i}$ starting from $j$-th joint when the robot is at the configuration $\boldsymbol{q}$. The term $\mathcal{R}_{i}$ designates the “incomplete” robot starting at its base and ending up at its $i$-th joint.